OpenStudy (anonymous):

If Zack makes 1½ quarts of lemonade, then he uses 6 lemons. If Zack makes 1½ quarts of lemonade, then he makes 4 servings. Zack uses 5 lemons. Which is a valid conclusion? A. Zack does not make 1½ quarts. B.Zack makes 3 servings. C.Zack makes 2 servings. D.Zack makes 1 quart.

OpenStudy (kropot72):

1 lemon makes \[\frac{1\frac{1}{2}}{6}\ quart\] 5 lemons make \[\frac{1\frac{1}{2}}{6}\times 5=\frac{3}{2}\times \frac{1}{6}\times\frac{5}{1}=?\ quarts\]